OpenStudy (anonymous):

The company discovers that it costs $103 to produce 3 widgets, $163 to produce 7 widgets, and $$328 to produce 12 widgets. Using the quadratic function, find the total cost of producing 2 widgets.

OpenStudy (amistre64):

this one?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i just put up the new question

OpenStudy (amistre64):

(3,103) (7,163) (12,328) ok...... -9a +-3b -c =-103 49a + 7b +c = 163 144a +12b +c = 328

OpenStudy (amistre64):

you want me to answer this one or that one?

OpenStudy (anonymous):

the other one. please!